### v1

# Particle Filtering for a Class of State-Space Models with Low and Degenerate Observational Noise

Abstract
We consider the discrete-time filtering problem in scenarios where the observation noise is degenerate or low. We focus on the case where the observation equation is a linear function of the state and that additive noise is low or degenerate, however, we place minimal assumptions on the hidden state process. In this scenario we derive new particle filtering (PF) algorithms and, under assumptions, in such a way that as the noise becomes more degenerate a PF which approximates the low noise filtering problem provably inherits the properties of the PF used in the degenerate case. We extend our framework to the case where the hidden states are drawn from a diffusion process. In this scenario we develop new PFs which are robust to both low noise and fine levels of time discretization. We illustrate our algorithms numerically on several examples.
Key Words: Filtering, particle filters, low and degenerate noise, Diffusions, time discretization.
Particle Filtering for a Class of State-Space Models with Low and Degenerate Observational Noise
BY ABYLAY ZHUMEKENOV1, ALEXANDROS BESKOS2, DAN CRISAN3, AJAY JASRA1 & NIKOLAS KANTAS3
1School of Data Science, The Chinese University of Hong Kong, Shenzen, Shenzhen, CN.
2Department of Statistical Science, University College London, London, WC1E 6BT, UK
3Department of Mathematics, Imperial College London, London, SW7 2AZ, UK
E-Mail:
abylayzhumekenov@cuhk.edu.cn; a.beskos@ucl.ac.uk;d.crisan@ic.ac.uk; ajayjasra@cuhk.edu.cn; n.kantas@ic.ac.uk
1 Introduction
The discrete-time filtering problem considers an unobserved Markov chain $\{X_{n}\}_{n\geq 0}$ (the signal or hidden chain) and associated data $\{Y_{n}\}_{n\geq 1}$. The data at any time $k$ are assumed to depend on $\{X_{n}\}_{n\geq 0}$ only by the position $X_{k}$ of the hidden state. The objective of filtering, assuming it is well-defined is to compute expectations with respect to (w.r.t.) the conditional distribution of the signal at each time $n$, given all the data observed up-to that time; we call such a distribution the filter. This problem has numerous applications in statistics, applied mathematics, finance and beyond; we refer the reader to the texts [2, 8, 9] and the references therein for a more complete summary.
In most cases of practical interest, the filter is unavailable analytically and one must resort to numerical approximation of which there are plethora of methods; see [2, 8, 9] for some coverage of the subject. In this article we focus on the approach called particle filters (PF). These are Monte Carlo, simulation-based methods that simulate a collection of $N\in\mathbb{N}$ samples in parallel and undergo what are called sampling and resampling operations. Such methods are very well understood from a mathematical perspective (e.g. [9]) and provide almost surely exact approximations of the filter as $N\rightarrow\infty$. Typically they are effective when the dimension of the hidden state, $d_{x}$ say, is moderate, although there are several attempts to enhance the methodology to some special-case high-dimensional problems (e.g. [14, 15]).
In this article we focus on the case where the observations have low or degenerate noise. For instance one could consider the case:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}$ (1.1)
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}}$, $\Delta\geq 0$, $\{\epsilon_{n}\}_{n\geq 1}$ a sequence of $\mathbb{R}^{d_{y}}$-valued independently and identically distributed (i.i.d.) noises and for technical reasons $1\leq d_{y}<d_{x}$, and $A_{n}$ a sequence of $d_{y}\times d_{x}$ real matrices of full row rank. The low noise case could be interpreted as $\Delta$ small (e.g. $\Delta=10^{-6}$) and the degenerate case with $\Delta=0$. As was noticed by [21], this scenario is important in many applications where it is expected that the signal dynamics is strongly driving the observed data; see [3, 12] for example. From a mathematical perspective, in the case $\Delta=0$ one has to define the filter on a sequence of manifolds $\mathsf{M}_{n}=\{x\in\mathbb{R}^{d_{x}}:y_{n}-A_{n}x=0\}$ and indeed this has been explored, for instance, in [21]. In the context of fitting a single model for example the smoothing distribution (that is of $X_{0},\dots,X_{n}|y_{1},\dots,y_{n}$ with $n$ fixed) sampling on manifolds has been quite well investigated using Markov chain Monte Carlo (MCMC) in several interesting contexts; see for example [5, 7, 10, 11, 19]. The extension to the filtering problem seems only to be considered in [20, 21], although [20] do not consider a manifold that has been defined by the data such as $\mathsf{M}_{n}$. The approach of [21] is to leverage sequential MCMC and as such relies on the afore-mentioned MCMC literature. The methodology there is particularly useful in high-dimensions ($d_{x}$ large) but can be computationally expensive for moderate dimensional problems that the PF itself is well-suited towards and it is this latter direction that we follow in this paper.
The application of PFs for low noise problems seldom works well: it is well-known and illustrated later in this article. In the worst case scenario the results can be very numerically unstable. The main idea of this article is to consider, even in the low noise case, sampling on a sequence of manifolds. For instance, in the context of (1.1) $\widetilde{\mathsf{M}}_{n}=\{(x,\epsilon)\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}:y_{n}-A_{n}x-\Delta^{1/2}\epsilon=0\}$. The idea then is if one can define a PF on $\mathsf{M}_{n}$ then the PF on $\widetilde{\mathsf{M}}_{n}$ will ultimately inherit the properties of the PF in the degenerate case ($\Delta=0$). In this paper we explore exactly this issue in the presence of linear observation models such as (1.1) and our contributions can be summarized as follows:
  • We derive new PFs for the low noise and degenerate case for linear observation models. We place minimal assumptions on the hidden process $\{X_{n}\}_{n\geq 0}$.
  • In the scenario of the previous bullet and under assumptions, we prove that the $\mathbb{L}_{r}$-error of the PF in the low noise case will converge as $\Delta\downarrow 0$ to the $\mathbb{L}_{r}$-error of the PF in the degenerate noise case.
  • For additive Gaussian errors in the hidden chain and observations we derive the optimal proposal and prove that the weights of the PF in the low noise case will converge as $\Delta\downarrow 0$ to the weights of the PF in the degenerate noise case.
  • In the case where the hidden dynamics follow a diffusion process, we derive a new PF that is robust not only to low noise but also to fine time discretization error. This appears to be the first PF of its kind.
The significance of this work is not only for filtering. The new PFs in this work could be used for parameter estimation associated to diffusion models, such as was done in [1, 13] using the multilevel Monte Carlo method. We note in terms of the fourth contribution, [6] develop a methodology for the case of hypoelliptic diffusions and degenerate noise (with linear constraints), stating that PFs can be used in their context. However, they do not make the point that this enables the low noise case, which is one of the main messages of this work.
This article is structured as follows. In Section 2 we present the collection of models we will consider. In Section 3 we give two particular types of models with PF algorithms and mathematical results. In Section 4 we show how to extend our methodology to the case where the hidden dynamics follow a diffusion process. In Section 5 we present our numerical results. The proofs of our mathematical results can be found in Appendix A.
2 Modeling
2.1 State-Space Model
We present a large class of state-space models in the following section. The main idea is to relate what we call the noise parameterization and degenerate noise case in full generality. As we stated in the introduction, however, we are only able to present computational methods for the case of linear observation regimes.
We consider the following state-space model (SSM) for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$ given:
$Y_{n}$ $=H_{n}(X_{n},\epsilon_{n})$
$X_{n}$ $=F_{n}(X_{n-1},\nu_{n})$
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}+d_{x}}$, $(d_{x},d_{y})\in\mathbb{N}^{2}$, $(\epsilon_{n},\nu_{n})\in\mathbb{R}^{d_{y}+d_{x}}$ are a sequence of independent noises (independent across $n$ and of each other) with positive and smooth Lebesgue density, $H_{n}:\mathbb{R}^{d_{x}+d_{y}}\rightarrow\mathbb{R}^{d_{y}}$, $F_{n}:\mathbb{R}^{2d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ are a sequence of continuously differentiable functions and whose Jacobian is of full row rank.
2.2 Natural Parametrization
In the given scenario we can define the filtering densities w.r.t. $d_{x}$-dimensional Lebesgue measure as follows. Denoting the positive conditional Lebesgue densities of $Y_{n}|X_{n}$ and $X_{n}|X_{n-1}$ as $h_{n}(y_{n}|x_{n})$ and $f_{n}(x_{n}|x_{n-1})$ respectively we have that the filter (see e.g. [2] and the references therein) at time 1 is
$\pi_{1}(x_{1})=\frac{h_{1}(y_{1}|x_{1})f_{1}(x_{1}|x_{0})}{\int_{\mathbb{R}^{d_{x}}}h_{1}(y_{1}|x_{1})f_{1}(x_{1}|x_{0})\lambda_{d_{x}}(dx_{1})}$
where $\lambda_{d_{x}}$ is $d_{x}$-dimensional Lebesgue measure and we supress the dependence of $y_{1}$ in the filter. The filter at any subsequent time point $n\geq 1$ has the following prediction-updating formula:
$\pi_{n}(x_{n})=\frac{h_{n}(y_{n}|x_{n})\int_{\mathbb{R}^{d_{x}}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}(x_{n-1})\lambda_{d_{x}}(dx_{n-1})}{\int_{\mathbb{R}^{d_{x}}}h_{n}(y_{n}|x_{1})\int_{\mathbb{R}^{d_{x}}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}(x_{n-1})\lambda_{d_{x}}(dx_{n-1})\lambda_{d_{x}}(dx_{n})}.$
In practice one cannot compute the filter and resorts to numerical methods. Consider when $d_{y}=1$ and $\epsilon_{n}\sim\mathcal{N}(0,\sigma_{n}^{2})$ with $\mathcal{N}(0,\sigma_{n}^{2})$ the one-dimensional normal distribution of mean zero and variance $\sigma_{n}^{2}$ very small say $10^{-6}$ we call this a special case of low observational noise. In such cases it can be difficult to design an efficient numerical methods and to help deal with such a scenario we will consider a SSM under a noise-based parameterization which we now detail.
2.3 Noise Parametrization and Degenerate Noise
Define the manifold $\widetilde{\mathsf{M}}_{n}=\{(x_{n},\epsilon_{n})\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{y}}:y_{n}-H_{n}(x_{n},\epsilon_{n})=0\}$. We set $c_{n}(x_{n},\epsilon_{n})=y_{n}-H_{n}(x_{n},\epsilon_{n})$. We shall suppose that $\mathbb{R}^{d_{x}}$ is equipped with a metric tensor with a fixed positive definite matrix representation $M$. Then it is simple to show that the filter density of $x_{n},\epsilon_{n}$ w.r.t. the Riemannian measure on $\widetilde{\mathsf{M}}_{n}$ denote $\mu^{M}_{\mathsf{M}_{n}}$ is as follows (see e.g. [21]). We set
$\widetilde{G}_{n}(x_{n},\epsilon_{n})=\textrm{det}(\partial c_{n}(x_{n},\epsilon_{n})M^{-1}\partial c_{n}(x_{n},\epsilon_{n})^{\top})^{-1/2}$
where $\partial c_{n}$ is the Jacobian matrix and det is the determinant. For time 1 we have
$\widetilde{\pi}_{1}(x_{1},\epsilon_{1})=\frac{\widetilde{G}_{1}(x_{1},\epsilon_{1})f(x_{1}|x_{0})p_{1}(\epsilon_{1})}{\int_{\widetilde{\mathsf{M}}_{1}}\widetilde{G}_{1}(x_{1},\epsilon_{1})f(x_{1}|x_{0})p_{1}(\epsilon_{1})\mu^{M}_{\widetilde{\mathsf{M}}_{1}}(d(x_{n},\epsilon_{n}))}.$
and at any subsequent time $n\geq 2$:
$\widetilde{\pi}_{n}(x_{n},\epsilon_{n})=\frac{\widetilde{G}_{n}(x_{n},\epsilon_{n})p_{n}(\epsilon_{n})\int_{\mathsf{M}_{n-1}}f(x_{n}|x_{n-1})\widetilde{\pi}_{n}(x_{n-1},\epsilon_{n-1})\mu^{M}_{\widetilde{\mathsf{M}}_{n-1}}(d(x_{n-1},\epsilon_{n-1}))}{\int_{\widetilde{\mathsf{M}}_{n}}\widetilde{G}_{n}(x_{n},\epsilon_{n})p(\epsilon_{n})\int_{\widetilde{\mathsf{M}}_{n-1}}f(x_{n}|x_{n-1})\widetilde{\pi}_{n}(x_{n-1},\epsilon_{n-1})\mu^{M}_{\widetilde{\mathsf{M}}_{n-1}}(d(x_{n-1},\epsilon_{n-1}))\mu^{M}_{\widetilde{\mathsf{M}}_{n}}(d(x_{n},\epsilon_{n}))}.$
Here $p_{n}(\epsilon_{n})$ is the Lebesgue density of $\epsilon_{n}$. We remark that one can work with the noise parameters of the hidden state as well. However, as we focus on observational noise, this representation will suffice. As we will establish in the next Section when the noise in the observation equation is low, one can consider numerical algorithms under the noise representation and ultimately, as the ‘noise disappears’ recover an algorithm associated to the degenerate noise filtering problem, the latter of which we now explain
2.3.1 Degenerate Noise
The approach here is easily adapted to the case of degenerate observation noise as we now explain. Writing the SSM as
$Y_{n}$ $=H_{n}^{\star}(X_{n})$
$X_{n}$ $=F_{n}(X_{n-1},\nu_{n})$
where $(Y_{n},X_{n})\in\mathbb{R}^{d_{y}+d_{x}}$, $(d_{x},d_{y})\in\mathbb{N}^{2}$, $d_{y}<d_{x}$, $\nu_{n}\in\mathbb{R}^{d_{x}}$ a sequence of independent noises (independent across $n$) with positive and smooth Lebesgue density, $H_{n}^{\star}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$, $F_{n}:\mathbb{R}^{2d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ are a sequence of continuously differentiable functions and whose Jacobian is of full row rank. Note that we require $d_{y}<d_{x}$ as we aim to work with an embedded manifold.
As above, define the manifold $\mathsf{M}_{n}^{\star}=\{x_{n}\in\mathbb{R}^{d_{x}}:y_{n}-H_{n}^{\star}(x_{n})=0\}$. We set $c_{n}(x_{n})=y_{n}-H_{n}^{\star}(x_{n})$. We shall suppose that $\mathbb{R}^{d_{x}}$ is equipped with a metric tensor with a fixed positive definite matrix representation $M$. We set
$G_{n}(x_{n})=\textrm{det}(\partial c_{n}(x_{n})M^{-1}\partial c_{n}(x_{n})^{\top})^{-1/2}.$
As before the filter densities w.r.t. $\mu_{M}^{\mathsf{M}_{n}^{\star}}$ can be written down as
$\pi_{1}^{\star}(x_{1})=\frac{G_{1}(x_{1})f_{1}(x_{1}|x_{0})}{\int_{\mathsf{M}_{1}^{\star}}G_{1}(x_{1})f_{1}(x_{1}|x_{0})\mu^{M}_{\mathsf{M}_{1}^{\star}}(dx_{n})}.$
and at any subsequent time $n\geq 2$:
$\pi_{n}^{\star}(x_{n})=\frac{G_{n}(x_{n})\int_{\mathsf{M}_{n-1}^{\star}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}^{\star}(x_{n-1})\mu^{M}_{\mathsf{M}_{n-1}^{\star}}(dx_{n-1})}{\int_{\mathsf{M}_{n}^{\star}}\tilde{G}_{n}(x_{n})\int_{\mathsf{M}_{n-1}^{\star}}f_{n}(x_{n}|x_{n-1})\pi_{n-1}^{\star}(x_{n-1})\mu^{M}_{\mathsf{M}_{n-1}^{\star}}(dx_{n-1})\mu^{M}_{\mathsf{M}_{n}^{\star}}(dx_{n})}.$
3 Particular Cases
In this Section we consider the case of general additive noise (Section 3.1) and derive PFs in the low and degenerate noise cases. We then examine (Section 3.2) the case where the hidden process has additive Gaussian noise along with the observation process as well and derive the optimal proposals for the low and degenerate noise models.
3.1 General Additive Noise
We consider the state-space model for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$ given:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}\\$
where, $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, $\epsilon_{n}\in\mathbb{R}^{d_{y}}$ are a sequence of independent noises with positive Lebesgue density $p_{n}$ and $\Delta\geq 0$. For the hidden dynamics, we do not make any constraints beyond the existence of a positive Lebesgue transition density which can be evaluated pointwise; it is written as $f_{n}(x_{n}|x_{n-1})$ as before. In order to write down a PF we will write the smoother (in the noise parameterization) w.r.t. Lebesgue measure as, due to the linear constraint, this is simple to do.
Remark 3.1.
In this article we present noise degeneracy as the case that in the observation equation, as $\Delta\downarrow 0$ there is no noise at all. However, one could consider, say $Y_{n}\in\mathbb{R}^{3}$ and an observation equation of the type
$Y_{n}=A_{n}X_{n}+\mathtt{P}(\Delta)\epsilon_{n}$
where we define
$\mathtt{P}(\Delta)=\begin{bmatrix}1&0&0\\ 1&1&0\\ \Delta&\Delta&\Delta\end{bmatrix}$.
In this case one could send $\Delta\downarrow 0$ which corresponds to removing the observation noise in the third observation component. This type of scenario could be dealt with in the forthcoming commentary, in the sense that one would define a filter (noise parameterization) and a corresponding algorithm in the low noise case and the case $\Delta=0$ and that there is some convergence of the low-noise case to this limiting and degenerate case. The main changes would be mainly notational and hence, for ease of exposition, we keep the simple scenario that we have started with.
3.1.1 Low Noise Case
We begin in the low noise case. Let $x_{n}^{\star}$ solve $y_{n}=A_{n}x$, $\text{ker}(A_{n})=\{x\in\mathbb{R}^{d_{x}}:A_{n}x=0\}$ and $V_{n}$ is the $d_{x}\times(d_{x}-d_{y})$ matrix whose columns are an orthonormal basis of $\text{ker}(A_{n})$. In this case we will be filtering a random variable $\widetilde{Z}_{n}\in\mathbb{R}^{d_{x}}$ and as it will prove useful below, we set $\widetilde{z}_{n}^{\top}=(z_{n}^{\top},\overline{z}_{n}^{\top})^{\top}$, with $(z_{n},\overline{z}_{n})\in\mathbb{R}^{d_{x}-d_{y}}\times\mathbb{R}^{d_{y}}$. Let $A_{n}^{\Delta}:=[A_{n},\Delta^{1/2}\mathtt{I}_{d_{y}}]$, with $\mathtt{I}_{d_{y}}$ the $d_{y}\times d_{y}$ identity matrix, be a $d_{y}\times(d_{x}+d_{y})$ matrix, define $\text{ker}(A_{n}^{\Delta}):=\{x\in\mathbb{R}^{d_{x}+d_{y}}:A_{n}^{\Delta}x=0\}$, let $V_{n}^{\Delta}$ be a $(d_{x}+d_{y})\times d_{x}$ matrix whose columns are an orthonormal basis of $\text{ker}(A_{n}^{\Delta})$ and $(\widetilde{x}_{n}^{\Delta})^{\top}=((x_{n}^{\star})^{\top},0^{\top})$ be a $(d_{x}+d_{y})$-dimensional vector that solves $y_{n}=A_{n}^{\Delta}\widetilde{x}_{n}^{\Delta}$. For any $z\in\mathbb{R}^{d_{x}}$ set
$u_{n}^{\Delta}(z)=\widetilde{x}_{n}^{\Delta}+V_{n}^{\Delta}z$
with $u_{n}^{\Delta}(z)^{\top}=(u_{n}^{\Delta}(z,x)^{\top},u_{n}^{\Delta}(z,\epsilon)^{\top})$, $u_{n}^{\Delta}(z,x)\in\mathbb{R}^{d_{x}}$, $u_{n}^{\Delta}(z,\epsilon)\in\mathbb{R}^{d_{y}}$ and we note that $y_{n}=A_{n}^{\Delta}u_{n}^{\Delta}(z)$.
Given this construction, at any time $n\geq 1$ we can write the joint smoothing density of $\widetilde{z}_{1:n}\in\mathbb{R}^{nd_{x}}$ w.r.t. $\bigotimes_{i=1}^{n}\lambda_{d_{x}}(d\widetilde{z}_{i})$ as
$\pi_{n}^{\Delta}(\widetilde{z}_{1:n})\propto\prod_{k=1}^{n}p_{k}(u_{k}^{\Delta}(\widetilde{z}_{k},\epsilon))f_{k}(u_{k}^{\Delta}(\widetilde{z}_{k},x)|u_{k-1}^{\Delta}(\widetilde{z}_{k-1},x))$
where we set $u_{0}^{\Delta}(z,x)=x_{0}$, for any $z\in\mathbb{R}^{d_{x}}$. We are assuming the probability density is well-defined for any $\Delta>0$, that is, the normalization is finite.
In such a scenario, one can use a standard particle filter, which we now detail. Let $\widetilde{q}_{n}^{\Delta}$ be a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}}$ and set for any $(z^{\prime},z)\in\mathbb{R}^{2d_{x}}$
$w_{n}^{\Delta}(z^{\prime},z)=\frac{p_{n}(u_{n}^{\Delta}(z,\epsilon))f_{n}(u_{n}^{\Delta}(z,x)|u_{n-1}^{\Delta}(z^{\prime},x))}{\widetilde{q}_{n}^{\Delta}(z|z^{\prime})}.$
Note that $\widetilde{q}_{n}^{\Delta}$ can depend on $\Delta$ and when $n=1$ we may (or may not) condition on $x_{0}$ in $\widetilde{q}_{n}^{\Delta}$ but write $\widetilde{q}_{n}^{\Delta}(\cdot|x_{0})$ from herein. Let $N\in\mathbb{N}$ be given and denote for any $n\geq 1$, $\widetilde{z}_{n}^{1:N}=(\widetilde{z}_{n}^{1}\dots,\widetilde{z}_{n}^{N})$ which represent the particles of the PF. It will prove useful to use the ancestor representation of the resampling indices so we will also be interested in $(\widetilde{a}_{n}^{1},\dots,\widetilde{a}_{n}^{N})\in\{1,\dots,N\}^{N}$ and we use the convention that $\widetilde{a}_{0}^{i}=i$ for each $i\in\{1,\dots,N\}$. We will also write $\widetilde{z}_{0}^{\widetilde{a}_{0}^{j}}=x_{0}$. The PF is then the following.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{1}^{i}|x_{0}$ i.i.d. from $\widetilde{q}_{1}^{\Delta}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $\widetilde{a}_{n-1}^{i}|\widetilde{z}_{n-2}^{1:N},\widetilde{z}_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{i}},\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\Delta}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\widetilde{z}_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{n}^{i}|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}}$ independently from $\widetilde{q}_{n}^{\Delta}(\cdot|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})$. Set $n=n+1$ and return to the start of the bullet point.
Let $\varphi:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}$ be a bounded and measurable function (write such functions as $\mathtt{B}_{b}(\mathbb{R}^{d_{x}})$) then we can estimate:
$\pi_{n}^{\Delta}(\varphi):=\int_{\mathbb{R}^{nd_{x}}}\varphi(\widetilde{z}_{n})\pi_{n}^{\Delta}(\widetilde{z}_{1:n})\bigotimes_{i=1}^{n}\lambda_{d_{x}}(d\widetilde{z}_{i})$
by using the estimator, obtained just before the end of the iterate step
$\pi_{n}^{\Delta,N}(\varphi):=\frac{\sum_{i=1}^{N}\varphi(\widetilde{z}_{n}^{i})w_{n}^{\Delta}(\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}},\widetilde{z}_{n}^{i})}{\sum_{i=1}^{N}w_{n}^{\Delta}(\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}},\widetilde{z}_{n}^{i})}.$
The convergence of this estimator and the choice of proposal has been extensively discussed in the literature; see for instance [2, 9] and the references therein. We remark that the optimal, in terms of minimizing the variance of the weights conditional on $\widetilde{z}_{n-1}$, is
$\widetilde{q}_{n}^{\Delta,\text{opt}}(\widetilde{z}_{n}|\widetilde{z}_{n-1})=\frac{p_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},\epsilon))f_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},x)|u_{n-1}^{\Delta}(\widetilde{z}_{n-1},x))}{\int_{\mathbb{R}^{d_{x}}}p_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},\epsilon))f_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},x)|u_{n-1}^{\Delta}(\widetilde{z}_{n-1},x))\lambda_{d_{x}}(d\widetilde{z}_{n})}.$
This can only be used in certain scenarios, one of which we investigate below.
3.1.2 Degenerate Noise Case
The degenerate case can be handled in a similar manner, which we now describe. For any $(n,z)\in\mathbb{N}\times\mathbb{R}^{d_{x}-d_{y}}$ set
$u_{n}^{\star}(z)=x_{n}^{\star}+V_{n}z$
where $V_{n}$ is the $d_{x}\times(d_{x}-d_{y})$ matrix whose columns form an orthonormal basis of $\text{ker}(A_{n})=\{x\in\mathbb{R}_{d_{x}}:A_{n}x=0\}$ and with the convention that $u_{0}^{\star}(z)=x_{0}$ for any $z\in\mathbb{R}^{d_{x}-d_{y}}$. At any time $n\geq 1$ we can write the joint smoothing density of $z_{1:n}\in\mathbb{R}^{n(d_{x}-d_{y})}$ w.r.t. $\bigotimes_{i=1}^{n}\lambda_{d_{x}-d_{y}}(dz_{i})$ as
$\pi_{n}^{\star}(z_{1:n})\propto\prod_{k=1}^{n}f_{k}(u_{k}^{\star}(z_{k})|u_{k-1}^{\star}(z_{k-1}))$
where we note that $u_{0}^{\star}(z_{0})=x_{0}$. Let $q_{n}^{\star}$ be a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}-d_{y}}$ and set for any $(z^{\prime},z)\in\mathbb{R}^{2(d_{x}-d_{y})}$
$w_{n}^{\star}(z^{\prime},z)=\frac{f_{n}(u_{n}^{\star}(z)|u_{n-1}^{\star}(z))}{q_{n}^{\star}(z|z^{\prime})}.$
When $n=1$ we may (or may not) condition on (some part of) $x_{0}$ in $q_{n}^{\star}$ but again write $q_{n}^{\star}(\cdot|x_{0})$ from herein. Denote for any $n\geq 1$, $z_{n}^{1:N}=(z_{n}^{1}\dots,z_{n}^{N})$, the particles of the PF. The ancestor variables are $(a_{n}^{1},\dots,a_{n}^{N})\in\{1,\dots,N\}^{N}$ and we use the convention that $a_{0}^{i}=i$ for each $i\in\{1,\dots,N\}$. We will also write $z_{0}^{a_{0}^{j}}=x_{0}$.
The PF is as follows.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $Z_{1}^{i}|x_{0}$ i.i.d. from $q_{1}^{\star}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}|z_{n-2}^{1:N},z_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{i}},z_{n-1}^{a_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\star}(z_{n-2}^{a_{n-2}^{s}},z_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $Z_{n}^{i}|z_{n-1}^{a_{n-1}^{i}}$ independently from $q_{n}^{\star}(\cdot|z_{n-1}^{a_{n-1}^{i}})$. Set $n=n+1$ and return to the start of the bullet point.
Let $\varphi\in\mathtt{B}_{b}(\mathbb{R}^{d_{x}-d_{y}})$ then we can estimate:
$\pi_{n}^{\star}(\varphi):=\int_{\mathbb{R}^{nd_{x}}}\varphi(z_{n})\pi_{n}^{\star}(z_{1:n})\bigotimes_{i=1}^{n}\lambda_{d_{x}-d_{y}}(dz_{i})$
by using the estimator, obtained just before the end of the iterate step
$\pi_{n}^{\star,N}(\varphi):=\frac{\sum_{i=1}^{N}\varphi(z_{n}^{i})w_{n}^{\star}(z_{n-1}^{a_{n-1}^{i}},z_{n}^{i})}{\sum_{i=1}^{N}w_{n}^{\star}(z_{n-1}^{a_{n-1}^{i}},z_{n}^{i})}.$
As in the low noise case, the optimal proposal here would be
$q_{n}^{\star,\text{opt}}(z_{n}|z_{n-1})=\frac{f_{n}(u_{n}^{\star}(z_{n})|u_{n-1}^{\star}(z_{n-1}))}{\int_{\mathbb{R}^{d_{x}-d_{y}}}f_{n}(u_{n}^{\star}(z_{n})|u_{n-1}^{\star}(z_{n-1}))\lambda_{d_{x}-d_{y}}(dz_{n})}.$
3.1.3 Mathematical Result
We now consider the convergence, in a sense to be made precise below, of the PF in the low noise case to that of the degenerate noise case. Recall that we use the notation for $\widetilde{z}\in\mathbb{R}^{d_{x}}$ as $\widetilde{z}^{\top}=(z^{\top},\overline{z}^{\top})$, with $z\in\mathbb{R}^{d_{x}}$. We make the following assumption.
  • (A1)
    For any $(n,\widetilde{z},\widetilde{z}^{\prime})\in\mathbb{N}\times\mathbb{R}^{2d_{x}}$ we have
    $\lim_{\Delta\downarrow 0}\widetilde{q}_{n}^{\Delta}(\widetilde{z}|\widetilde{z}^{\prime})=p_{n}(\overline{z})q_{n}^{\star}(z|z^{\prime}).$
This assumption is not overly restrictive as one chooses the proposal densities. The assumption means that as the noise disappears, one should propose the $d_{x}-d_{y}$ random variable, just as in the in the degenerate noise case and the noise variable according to the noise density.
We denote by $\mathbb{E}^{\Delta}$ as expectations w.r.t. the law of the PF in the low noise case and by $\mathbb{E}^{\star}$ as expectations w.r.t. the law of the PF in the degenerate noise case. We can now state our main result whose proof is in Appendix A.1.
Proposition 3.1.
Assume (A(A1)). Then for any $(n,r,\varphi)\in\mathbb{N}^{2}\times\mathtt{B}_{b}(\mathbb{R}^{d_{x}-d_{y}})$ we have
$\lim_{\Delta\downarrow 0}\mathbb{E}^{\Delta}\left[\left|\pi_{n}^{\Delta,N}(\varphi)-\pi_{n}^{\Delta}(\varphi)\right|^{r}\right]^{1/r}=\mathbb{E}^{\star}\left[\left|\pi_{n}^{\star,N}(\varphi)-\pi_{n}^{\star}(\varphi)\right|^{r}\right]^{1/r}.$
Remark 3.2.
The main purpose of Proposition 3.1 is to verify that the properties of the low noise case PF are inherited, as $\Delta\downarrow 0$ to that of the degenerate noise case. In other words if one can design a good PF for the degenerate noise case, then the low noise filtering problem can be effectively handled. For instance if one has that
$\sup_{n\geq 1}\sup_{(z^{\prime}z)\in\mathbb{R}^{2(d_{x}-d_{y})}}w_{n}^{\star}(z^{\prime},z)<+\infty$
and
$\inf_{n\geq 1}\inf_{(z^{\prime}z)\in\mathbb{R}^{2(d_{x}-d_{y})}}q_{n}(z|z^{\prime})$ $>0$
$\sup_{n\geq 1}\sup_{(z^{\prime}z)\in\mathbb{R}^{2(d_{x}-d_{y})}}q_{n}(z|z^{\prime})$ $<+\infty$
then by standard results for PFs (e.g. [9]) we have that
$\sup_{n\geq 1}\mathbb{E}^{\star}\left[\left|\pi_{n}^{\star,N}(\varphi)-\pi_{n}^{\star}(\varphi)\right|^{r}\right]^{1/r}=\mathcal{O}(N^{-\tfrac{1}{2}})$
and so the time uniform errors of the degenerate noise PF are then then obtained for the low noise PF.
3.2 Special Case of Additive Noise
We consider the following state-space model for $n\in\mathbb{N}$ and $X_{0}\in\mathbb{R}^{d_{x}}$ given:
$Y_{n}$ $=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}$
$X_{n}$ $=f_{n}(X_{n-1})+\nu_{n}$
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, $\Delta\geq 0$ $f_{n}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$. Independently for each $n\in\mathbb{N}$ $\epsilon_{n}\sim\mathcal{N}_{d_{y}}(0,\Sigma_{n})$ and independently $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega_{n})$ where, $\Sigma_{n},\Omega_{n}$ are a sequence of real positive definite symmetric matrices and for $d\in\mathbb{N}$ fixed, $(\mu,\kappa)\in\mathbb{R}^{d}\times\mathbb{R}^{d\times d}$ ($\kappa$ is also positive definite and symmetric) $\mathcal{N}_{d}(\mu,\kappa)$ is the $d$-dimensional Gaussian distribution with mean $\mu$ and covariance $\kappa$.
3.2.1 Particle Filter for the Natural Parameterization
For this problem we already know the PF for the filter in the noise parameterization case as well as the degenerate noise case. For completeness, we detail the natural parameterization case when $\Delta>0$. $q_{n}$ is a sequence of positive Markov transition densities on $\mathbb{R}^{d_{x}}$ and we set for any $(x^{\prime},x)\in\mathbb{R}^{2d_{x}}$
$w_{n}(x^{\prime},x)=\frac{h_{n}(y_{n}|x)f_{n}(x|x^{\prime})}{q_{n}(x|x^{\prime})}.$
As before we set $a_{0}^{i}=i$ for $i\in\{1,\dots,n\}$ and $x_{0}^{i}=x_{0}$. The PF is then the following.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $X_{1}^{i}|x_{0}$ i.i.d. from $q_{1}(\cdot|x_{0})$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}|x_{n-2}^{1:N},x_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}(x_{n-2}^{a_{n-2}^{i}},x_{n-1}^{a_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}(x_{n-2}^{a_{n-2}^{s}},x_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $X_{n}^{i}|x_{n-1}^{a_{n-1}^{i}}$ independently from $q_{n}(\cdot|x_{n-1}^{a_{n-1}^{i}})$. Set $n=n+1$ and return to the start of the bullet point.
The optimal proposal in this case is
$q_{n}^{\text{opt}}(x_{n}|x_{n-1})=\frac{h_{n}(y_{n}|x_{n})f_{n}(x_{n}|x_{n-1})}{\int_{\mathbb{R}^{d_{x}}}h_{n}(y_{n}|x_{n})f_{n}(x_{n}|x_{n-1})\lambda_{d_{x}}(dx_{n})}.$
Often used (and maligned) is the bootstrap PF where one uses $q_{n}(x|x^{\prime})=f_{n}(x|x^{\prime})$.
3.2.2 Mathematical Result
We now consider the optimal or bootstap proposals and the importance weights in the natural parameterization, noise parameterization (low noise) and the degenerate noise case. To aid the notations in the below statement we partition the matrix $V_{n}^{\Delta}$ into its first $d_{x}$ rows denoted by the $d_{x}\times d_{x}$ matrix $V_{n}^{\Delta}(x)$ and its last $d_{y}$ rows denoted by the $d_{y}\times d_{x}$ matrix $V_{n}^{\Delta}(\epsilon)$. We will also use the notation $x_{0}^{\star}+V^{\Delta}_{0}(x)z=x_{0}$ $x_{0}^{\star}+V_{0}z=x_{0}$ and where we have not defined the objects on L.H.S.. We have the following result whose proof is in Appendix A.2.
Proposition 3.2.
Consider particle filters associated to the model at the start of Section 3.2.
  1. 1.
    Natural parameterization with bootstrap proposal:
    $X_{n}|x_{n-1}$ $\sim\mathcal{N}_{d_{x}}(f(x_{n}),\Omega_{n})$
    $w_{n}(x_{n})$ $=\exp\left\{-\frac{1}{2\Delta}(y_{n}-A_{n}x_{n})^{\top}\Sigma^{-1}(y_{n}-A_{n}x_{n})\right\}$
  2. 2.
    Low noise with optimal proposal:
    $\widetilde{Z}_{n}|\widetilde{z}_{n-1}$ $\sim\mathcal{N}_{d_{x}}(\widetilde{\mu}_{n},\widetilde{\Omega}_{n})$
    $w_{n}^{\Delta,\text{\emph{opt}}}(\widetilde{z}_{n-1})$ $=\exp\left\{\frac{1}{2}\left(\widetilde{\mu}_{n}^{\top}\widetilde{\Omega}_{n}^{-1}\widetilde{\mu}_{n}-\widetilde{\mathtt{m}}_{n}^{\top}\Omega_{n}^{-1}\widetilde{\mathtt{m}}_{n}\right)\right\}$
    where
    $\widetilde{\mu}_{n}$ $=\widetilde{\Omega}_{n}V_{n}^{\Delta}(x)^{\top}\Omega_{n}^{-1}(f_{n}(x_{n-1}^{\star}+V_{n-1}^{\Delta}(x)z_{n-1})-x_{n}^{\star})$
    $\widetilde{\Omega}_{n}^{-1}$ $=V_{n}^{\Delta}(\epsilon)^{\top}\Sigma_{n}^{-1}V_{n}^{\Delta}(\epsilon)+V_{n}^{\Delta}(x)^{\top}\Omega_{n}^{-1}V_{n}^{\Delta}(x)$
    $\widetilde{\mathtt{m}}_{n}$ $=f_{n}(x_{n-1}^{\star}+V_{n-1}^{\Delta}(x)z_{n-1})-x_{n}^{\star}.$
  3. 3.
    Degenerate noise with optimal proposal:
    $Z_{n}|z_{n-1}$ $\sim\mathcal{N}_{d_{x}-d_{y}}(\mu_{n}^{\star},\Omega_{n}^{\star})$
    $w_{n}^{\star,\text{\emph{opt}}}(z_{n-1})$ $=\exp\left\{\frac{1}{2}\left((\mu_{n}^{\star})^{\top}(\Omega_{n}^{\star})^{-1}\mu_{n}^{\star}-(\mathtt{m}_{n}^{\star})^{\top}\Omega_{n}^{-1}\mathtt{m}_{n}^{\star}\right)\right\}$
    where
    $\mu_{n}^{\star}$ $=\Omega_{n}^{\star}V_{n}^{\top}\Omega_{n}^{-1}(f_{n}(x_{n-1}^{\star}+V_{n-1}z_{n-1})-x_{n}^{\star})$
    $(\Omega_{n}^{\star})^{-1}$ $=V_{n}^{\top}\Omega_{n}^{-1}V_{n}$
    $\mathtt{m}_{n}^{\star}$ $=f_{n}(x_{n-1}^{\star}+V_{n-1}z_{n-1})-x_{n}^{\star}.$
Finally, we have
$\lim_{\Delta\downarrow 0}w_{n}^{\Delta,\text{\emph{opt}}}(\widetilde{z}_{n-1})=w_{n}^{\star,\text{\emph{opt}}}(z_{n-1}).$
Remark 3.3.
The significance of Proposition 3.2 is as follows. $\Delta$ controls the amount of noise and the low noise case is as $\Delta\downarrow 0$. In the case of the bootstrap PF it is clear that numerically that $w_{n}(x_{n})$ will degenerate to zero and hence this algorithm is of little use; this is unsuprising as the proposal does not take into account any information associated to the data. We remark that for the optimal proposal in the case of the natural parameterization the weight does not collapse or explode, but clearly the proposal and target (which begins to concentrate on $\mathsf{M}_{n}$) will become mutually singular as $\Delta\downarrow 0$ so it is of little use in the very small noise and degenerate case. In the low and degenerate noise cases (2 & 3 in Proposition 3.2) one can characterize the optimal proposal and the associated weight and, as was the case in Proposition 3.1, there is a convergence of the weights from the low noise to degenerate case with a similar conclusion to Remark 3.2.
4 Partially Observed Diffusions
4.1 Diffusion Process
We will now consider the problem of filtering diffusion processes in the low and degenerate noise case. We consider the following diffusion process on the filtered probability space $(\Omega,\mathscr{F},\{\mathscr{F}_{t}\}_{t\geq 0},\mathbb{P})$.
$dX_{t}=\mu(X_{t})dt+\sigma(X_{t})dW_{t}\quad\quad X_{0}=x_{0}$ (4.1)
where $X_{t}\in\mathbb{R}^{d_{x}}$, $\mu:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$, and $\sigma:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}\times d_{x}}$ and $\{W_{t}\}_{t\geq 0}$ is a standard $d_{x}$-dimensional Brownian motion. Denote $\mu^{j}$ as the $j^{th}$-component of $\mu$, $j\in\{1,\dots,d_{x}\}$ and $\sigma^{j,k}$ as the $(j,k)^{th}$-component of $\sigma$, $(j,k)\in\{1,\dots,d_{x}\}^{2}$. As the exposition will closely follow the works in [1, 4, 13, 17] we make the following assumption which is also made in most of those papers. Below $\mathcal{C}(\mathbb{R}^{d_{x}})$ are the twice continuously differentiable functions on $\mathbb{R}^{d_{x}}$ and $\|\cdot\|$ is the $L_{2}$-norm.
We have $\mu^{j}\in\mathcal{C}^{2}(\mathbb{R}^{d_{x}})$ and $\sigma^{j,k}\in\mathcal{C}^{2}(\mathbb{R}^{d_{x}})$, for $(j,k)\in\{1,\ldots,d\}^{2}$. In addition, $\mu$ and $\sigma$ satisfy
  • (i)
    uniform ellipticity: $\Sigma(x):=\sigma(x)\sigma(x)^{\top}$ is uniformly positive definite over $x\in\mathbb{R}^{d_{x}}$;
  • (ii)
    globally Lipschitz: There exists a constant $C<\infty$ such that $|\mu^{j}(x)-\mu^{j}(x^{\prime})|+|\sigma^{j,k}(x)-\sigma^{j,k}(x^{\prime})|\leq C\|x-x^{\prime}\|$ for all $(x,x^{\prime})\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}$, $(j,k)\in\{1,\dots,d_{x}\}^{2}$.
We note that one could use the Euler-Maruyama time-discretization method to obtain the transition density of the hidden process. However, it is well-known (e.g. [4] and the references therein) that this method, when combined with a PF, typically collapses (in some sense) as the time-discretization becomes finer and thus we consider a bridge construction as often used in the literature. As a result the PFs that we will present in Sections 4.3.1 and 4.3.2 will be robust to the case of low noise and fine time discretizations.
4.2 Bridge Construction
We review the method in [16, 18] as has been used in [4]. The exposition closely follows that in [1, 4, 13]. Consider the case $t\in[s_{1},s_{2}]$, $0\leq s_{1},s_{2}\leq T$ and let $\mathbf{X}_{[s_{1},s_{2}]}:=\{X_{t}\}_{t\in[s_{1},s_{2}]}$, and $\mathbf{W}_{[s_{1},s_{2}]}:=\{W_{t}\}_{t\in[s_{1},s_{2}]}$. Set $f_{t,s_{2}}(x^{\prime}|x)$ denote the unknown transition density from time $t$ to $s_{2}$ associated to (4.1). If one could sample from $f_{s_{1},s_{2}}$ to obtain $(x,x^{\prime})\in\mathbb{R}^{2d_{x}}$, we will explain that we can interpolate these points by using a bridge process. This process will have a drift given by $\mu(x)+\Sigma(x)\nabla_{x}\log{f}_{t,s_{2}}(x^{\prime}|x)$. Let $\mathbb{P}_{x,x^{\prime}}$ denote the law of the solution of the SDE (4.1), on $s_{1},s_{2}]$, started at $x$ and conditioned to hit $x^{\prime}$ at time $s_{2}$.
We introduce a user-specified auxiliary process $\{\tilde{X}_{t}\}_{t\in[s_{1},s_{2}]}$ following:
$d\tilde{X}_{t}=\tilde{\mu}(t,\tilde{X}_{t})dt+\tilde{\sigma}(t,\tilde{X}_{t})dW_{t},\quad t\in[s_{1},s_{2}],\quad\penalty 10000\ \tilde{X}_{s_{1}}=x,$ (4.2)
where $\tilde{\mu}:[s_{1},s_{2}]\times\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ and $\tilde{\sigma}:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}\times d_{x}}$ is such that
$\tilde{\Sigma}(s_{2},x^{\prime}):=\tilde{\sigma}(s_{2},x^{\prime})\tilde{\sigma}(s_{2},x^{\prime})^{\top}\equiv\Sigma(x^{\prime}).$
(4.2) is specified so that its transition density $\tilde{f}$ is available; see [16, Section 2.2] for the technical conditions on $\tilde{\mu},\tilde{\Sigma},\tilde{f}$. The main purpose of $\{\tilde{X}_{t}\}_{t\in[s_{1},s_{2}]}$ is to sample $x^{\prime}$ and use its transition density to construct another process $\{X_{t}^{\circ}\}_{t\in[s_{1},s_{2}]}$ conditioned to hit $x^{\prime}$ at $t=s_{2}$; which in turn will be an importance proposal for $\{X_{t}\}_{t\in[s_{1},s_{2}]}$. Set:
$dX^{\circ}_{t}=\mu_{s_{2}}^{\circ}(t,X^{\circ}_{t};x^{\prime})dt+\sigma(X^{\circ}_{t})dW_{t},\quad t\in[s_{1},s_{2}],\quad\penalty 10000\ X^{\circ}_{s_{1}}=x,$ (4.3)
where:
$\mu_{s_{2}}^{\circ}(t,x;x^{\prime})=\mu(x)+\Sigma(x)\nabla_{x}\log\tilde{f}_{t,s_{2}}(x^{\prime}|x),$
and denote by $\mathbb{P}^{\circ}_{x,x^{\prime}}$ the probability law of the solution of (4.3). The SDE in (4.3) yields:
$\mathbf{W}\rightarrow C_{s_{1},s_{2}}(x,\mathbf{W}_{[s_{1},s_{2}]},x^{\prime}),$ (4.4)
mapping the driving Wiener noise $\mathbf{W}$ to the solution of (4.3), reparametering the problem from $\mathbf{X}$ to $(\mathbf{W},x^{\prime})$.
[16] prove that $\mathbb{P}_{x,x^{\prime}}$ and $\mathbb{P}^{\circ}_{x,x^{\prime}}$ are absolutely continuous w.r.t. each other, with Radon-Nikodym derivative:
$\frac{d\mathbb{P}_{x,x^{\prime}}}{d\mathbb{P}^{\circ}_{x,x^{\prime}}}(\mathbf{X}_{[s_{1},s_{2}]})=\exp\Big\{\int_{s_{1}}^{s_{2}}L_{s_{2}}(t,X_{t})dt\Big\}\times\frac{\tilde{f}_{s_{1},s_{2}}(x^{\prime}|x)}{f_{s_{1},s_{2}}(x^{\prime}|x)},$ (4.5)
where:
$L_{s_{2}}$ $(t,x):=\left(\mu(x)-\tilde{\mu}(t,x)\right)^{\top}\,\nabla_{x}\log\tilde{f}_{t,s_{2}}(x^{\prime}|x)$
$-\frac{1}{2}\textrm{Tr}\,\Big\{\,\big[\Sigma(x)-\tilde{\Sigma}(t,x)\big]\big[-\nabla_{x}^{2}\log\tilde{f}_{t,s_{2}}(x^{\prime}|x)-\nabla_{x}\log\tilde{f}_{t,s_{2}}(x^{\prime}|x)\nabla_{x}\log\tilde{f}_{t,s_{2}}(x^{\prime}|x)^{\top}\big]\,\Big\},$
with $\textrm{Tr}(\cdot)$ denoting the trace of a squared matrix. Set
$R_{s_{1},s_{2}}(\mathbf{X}_{[s_{1},s_{2}]}):=\frac{d\mathbb{P}_{x,x^{\prime}}}{d\mathbb{P}^{\circ}_{x,x^{\prime}}}(\mathbf{X}_{[s_{1},s_{2}]})f_{s_{1},s_{2}}(x^{\prime}|x).$
4.3 Bridges and Manifolds
We will consider a linear observation process:
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}$
where $Y_{n}\in\mathbb{R}^{d_{y}}$, $1\leq d_{y}<d_{x}$, $A_{n}\in\mathbb{R}^{d_{y}\times d_{x}}$ is of full row rank, and independently for each $n\in\mathbb{N}$ $\epsilon_{n}\in\mathbb{R}^{d_{y}}$ are a sequence of independent noises with positive Lebesgue density $p_{n}$ and $\Delta\geq 0$ and $\Sigma_{n}$ is a sequence of real positive definite symmetric matrices. This is the same linear set-up as the previous section. Note that we present the model with observations at regular and unit times only to simplify the notation and make it consistent with previous articles [4]. However, this need not be the case and observations could in principle have a different amount of time in-between them, or indeed be irregularly observed.
4.3.1 Low Noise
Let $\mathbb{W}$ is the law of Brownian motion. The probability measure associated to the smoother in continuous-time and up-to a constant is at time $n$
$\prod_{k=1}^{n}\Big\{R_{k-1,k}(C_{k-1,k}(u_{k-1}^{\Delta}(\widetilde{z}_{k-1},x),\mathbf{w}_{[k-1,k]},u_{k}^{\Delta}(\widetilde{z}_{k},x)))\mathbb{W}(d\mathbf{w}_{[k-1,k]})p_{k}(u_{k}^{\Delta}(\widetilde{z}_{k},\epsilon))\lambda_{d_{x}}(d\widetilde{z}_{k})\Big\}.$ (4.6)
In practice we cannot work with this and consider a time discretization.
We first give the standard Euler-Maruyama time discretization of the solution to (4.3) (associated to a time interval [k1,k][k-1,k], $k\in\{1,\dots\}$) on a regular grid of spacing $\Xi_{l}=2^{-l}$, with starting point $x_{k-1}^{\circ}$ and ending point $x_{k}^{\circ}$. The choice $\Xi_{l}=2^{-l}$, is again for consistency of presentation with [4], but other schemes can be used in practice. We have, for $j\in\{0,1\dots,\Xi_{l}^{-1}-2\}$:
$X_{k-1+(j+1)\Delta_{l}}^{\circ}$ $=X_{k-1+j\Xi_{l}}^{\circ}+\mu_{k}^{\circ}(k-1+j\Xi_{l},X^{\circ}_{k-1+j\Xi_{l}};x_{k}^{\circ})\Xi_{l}+$
$\sigma(X^{\circ}_{k-1+j\Xi_{l}})\left[W_{k-1+(j+1)\Xi_{l}}-W_{k-1+j\Xi_{l}}\right].$ (4.7)
Given $(x_{k-1}^{\circ},x_{k}^{\circ})$ and $\mathbf{W}_{[k-1,k]}^{l}=(W_{k-1+\Xi_{l}}-W_{k-1},\dots,W_{k-\Xi_{l}}-W_{k-2\Xi_{l}})$, the recursion (4.3.1) induces a discretized path $X_{k-1+\Xi_{l}}^{\circ},\dots,X_{k-\Xi_{l}}^{\circ}$ and we write such a path, including the starting and ending points with the notation
$C_{k-1,k}^{l}(x_{k-1}^{\circ},\mathbf{W}_{[k-1,k]}^{l},x_{k}^{\circ}).$
We also need a discretization of the Radon-Nikodym derivative. Consider
$\mathbf{X}_{[k-1,k]}^{l}=(X_{k-1},X_{k-1+\Xi_{l}},\dots,X_{k})$
then we set
$R_{k-1,k}^{l}(\mathbf{X}_{[k-1,k]}^{l}):=\exp\Big\{\sum_{j=0}^{\Xi_{l}^{-1}-1}L_{k}(t,X_{k-1+j\Xi_{l}})\Xi_{l}\Big\}\tilde{f}_{k-1,k}(X_{k}|X_{k-1}).$
As a result of the time-discretization we focus on the smoother (up-to a normalizing constant):
$\prod_{k=1}^{n}\Big\{R_{k-1,k}^{l}(C_{k-1,k}^{l}(u_{k-1}^{\Delta}(\widetilde{z}_{k-1},x),\mathbf{w}_{[k-1,k]}^{l},u_{k}^{\Delta}(\widetilde{z}_{k},x)))\mathbb{W}(d\mathbf{w}_{[k-1,k]}^{l})p_{k}(u_{k}^{\Delta}(\widetilde{z}_{k},\epsilon))\lambda_{d_{x}}(d\widetilde{z}_{k})\Big\}$
We set
$w_{n}^{\Delta,l}(\widetilde{z}_{n-1},\mathbf{w}_{[n-1,n]}^{l},\widetilde{z}_{n})$ $=R_{n-1,n}^{l}(C_{n-1,n}^{l}(u_{n-1}^{\Delta}(\widetilde{z}_{n-1},x),\mathbf{w}_{[n-1,n]}^{l},u_{n}^{\Delta}(\widetilde{z}_{n},x)))\frac{p_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},\epsilon))}{\widetilde{q}_{n}^{\Delta}(\widetilde{z}_{n}|\widetilde{z}_{n-1})}.$
We can now modify the PF of Section 3.1.1 to this case as follows.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{1}^{i}|x_{0}$ i.i.d. from $\widetilde{q}_{1}^{\Delta}(\cdot|x_{0})$. and independently $\mathbf{w}_{[0,1]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $\widetilde{a}_{n-1}^{i}|\widetilde{z}_{n-2}^{1:N},\mathbf{w}_{[n-2,n-1]}^{1:N,l},\widetilde{z}_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{i}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{i},l},\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\Delta,l}(\widetilde{z}_{n-2}^{\widetilde{a}_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{\widetilde{a}_{n-2}^{s},l},\widetilde{z}_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $\widetilde{Z}_{n}^{i}|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}}$ independently from $\widetilde{q}_{n}^{\Delta}(\cdot|\widetilde{z}_{n-1}^{\widetilde{a}_{n-1}^{i}})$ and independently $\mathbf{w}_{[n-1,n]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=n+1$ and return to the start of the bullet point.
4.3.2 Degenerate Noise
In the degenerate noise case, one can follow the same path as the low noise case and write the time discretized smoother up-to a normalizing constant as
$\prod_{k=1}^{n}\Big\{R_{k-1,k}^{l}(C_{k-1,k}^{l}(u_{k-1}^{\star}(z_{k-1}),\mathbf{w}_{[k-1,k]}^{l},u_{k}^{\star}(z_{k})))\mathbb{W}(d\mathbf{w}_{[k-1,k]}^{l})\lambda_{d_{x}-d_{y}}(dz_{k})\Big\}.$
We set
$w_{n}^{\star,l}(z_{n-1},\mathbf{w}_{[n-1,n]}^{l},z_{n})$ $=R_{n-1,n}^{l}(C_{n-1,n}^{l}(u_{n-1}^{\star}(z_{n-1}),\mathbf{w}_{[n-1,n]}^{l},u_{n}^{\star}(z_{n})))\frac{1}{q_{n}^{\star}(z_{n}|z_{n-1})}.$
We can now modify the PF of Section 3.1.2 to this case as follows.
  • Initialize: For $i\in\{1,\dots,N\}$, sample $Z_{1}^{i}|x_{0}$ i.i.d. from $q_{1}^{\star}(\cdot|x_{0})$ and independently $\mathbf{w}_{[0,1]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=2$.
  • Iterate: For $i\in\{1,\dots,N\}$, sample $a_{n-1}^{i}|z_{n-2}^{1:N},\mathbf{w}_{[n-2,n-1]}^{1:N,l},z_{n-1}^{1:N}$ i.i.d. from
    $\frac{w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{i}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{i}},z_{n-1}^{a_{n-1}^{i}})}{\sum_{s=1}^{N}w_{n-1}^{\star,l}(z_{n-2}^{a_{n-2}^{s}},\mathbf{w}_{[n-2,n-1]}^{a_{n-2}^{s}},z_{n-1}^{s})}.$
    For $i\in\{1,\dots,N\}$, sample $Z_{n}^{i}|z_{n-1}^{a_{n-1}^{i}}$ independently from $q_{n}^{\star}(\cdot|z_{n-1}^{a_{n-1}^{i}})$ and independently $\mathbf{w}_{[n-1,n]}^{i,l}\sim\mathbb{W}(\cdot)$. Set $n=n+1$ and return to the start of the bullet point.
The algorithm here is presumably the uniformly elliptic embodyment of the ideas in [6].
5 Numerical Results
For all numerical examples we have
$Y_{n}=A_{n}X_{n}+\Delta^{1/2}\epsilon_{n}$ (5.1)
where, independently for each $n\in\mathbb{N}$, $\epsilon_{n}\sim\mathcal{N}_{d_{y}}(0,\mathtt{I}_{d_{y}})$ and varying $\Delta$. In Section 5.1 $d_{y}=1$, in Section 5.2 $d_{y}=2$ and in Section 5.3 $d_{y}=1$.
5.1 Linear Gaussian Model
We consider a linear Gaussian model with $d_{x}=10$
$X_{n}=B_{n}X_{n-1}+\nu_{n}$ (5.2)
where, independently for each $n\in\mathbb{N}$, $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega)$. We choose $X_{0}=(0,0)^{\top}$, $\Omega=\mathtt{I}_{d_{x}}$, $B_{n}=0.9\mathtt{I}_{d_{x}}$, $(A_{n})_{ij}=1/d_{x}$ (computes the average of the state), $n\in\{1,\dots,20\}$ and $\Delta=10^{-4}$. We run the Kalman filter, the particle filters with bootstrap and optimal proposals under the natural parameterization and the proposed filter under the noise parameterization. We set the number of particles to $N=10^{4}$ and resample if the effective sample size (ESS) drops below $N/2$. In Figure 1, we plot the marginal densities of the first component of $X_{n}$ at various $n$. We see that, except the bootstrap, all filters yield similar estimates as one might expect in this case.
Figure 1: Marginals densities of $X_{n,1}$ for $n\in\{1,6,11,16\}$. This is for the linear Gaussian model.
We now repeat the experiment for various $\Delta$ and alllow $\Delta\to 0$. We monitor the median ESS (Figure 2) across the time-steps for a given $\Delta$. In Figure 2 we observe that the bootstrap PF degenerates for relatively large $\Delta$, whilst the other methods remain stable. The optimal proposal under natural parameterization becomes unstable for very tiny $\Delta$, due to numerical errors and the reasoning in Remark 3.3. We can also see it in the mean squared error (MSE) of the mean estimates (error averaged across time), where the bootstrap PF collapses but the low-noise optimal proposal converges to the degenerate case.
Figure 2: ESS (left) and MSE (right) versus $\log_{10}(1/\Delta)$. This is for the linear Gaussian model.
5.2 Nonlinear Model with Additive Noise
Now we consider a nonlinear Lorenz-96 model with additive Gaussian noise, $d_{x}=8$
$X_{n}=F(X_{n-1})+\nu_{n}$ (5.3)
where $F_{i}(x)=x_{i}+\Delta t((x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F_{0})$, $i\in\{1,\dots,d_{x}\}$ and, independently for each $n\in\mathbb{N}$, $\nu_{n}\sim\mathcal{N}_{d_{x}}(0,\Omega)$. We fix the parameter of the model to $F_{0}=8$ to make the dynamics chaotic, and simulate it at stepsize $\Delta t=10^{-2}$ for $n\in\{1,\dots,400\}$ to obtain the data, starting from $X_{0}=(1,1)^{\top}$. We fix $\Delta=10^{-4}$ and $\Omega=\Delta t\mathtt{I}_{d_{x}}$, the matrix $A_{n}$ is constructed such that we observe the $1^{\text{s}t}$ and $5^{\text{t}h}$ component of $X_{n}$. We can observe the chaotic behavior of the model in Figure 3, which are the simulated dynamics of the model.
Figure 3: Components of $X_{n}$ for $n\in\{1,\dots,400\}$. This is for the nonlinear model.
We run all the mentioned filters (except for Kalman filter which does not apply here) again for $N=10^{4}$ particles with the same resampling rule. We plot several marginals in Figure 4, where we recover similar results as in the case of the linear Gaussian model, except for the case of the bootstrap PF; the algoritm performs so badly that the plots are essentially useless. Further, we plot the ESS versus the time index in Figure 5, where we see that the bootstrap PF collapses from the very beginning, while other methods give adequate results in terms of this measure. To further examine the accuracy of the PFs we observe individual sample paths of the partices (for a given run) and their mean for $8^{\text{t}h}$ component in Figure 6. In both the top and bottom of Figure 6 the true value is the actual signal used to generate the data. Apart from the bootstrap PF all of the methods are seemingly able to recover the true hidden state in the $8^{\text{t}h}$ component. In Figure 7 we consider re-running all of the methods for different $\Delta$ and sending $\Delta\to 0$. Figure 7 displays the median ESS (accross time) of the four PFs. In this case the optimal proposal under the natural parameterization does not seem to collapse (in terms of ESS) but this could be as the weight does not degenerate to zero.
Figure 4: Marginals densities of $X_{n,8}$ for $n\in\{1,101,201,301\}$. This is for the nonlinear model.
Figure 5: ESS versus $n$. This is for the nonlinear model.
Figure 6: Sample paths (top) and mean of the filter (bottom) for $X_{n,8}$. This is for the nonlinear model.
Figure 7: ESS versus $\log_{10}(1/\Delta)$. This is for the nonlinear model.
5.3 Diffusion Model
We now consider the FitzHugh-Nagumo diffusion model with $d_{x}=2$:
$dX_{t}=\mu(X_{t})dt+\sigma(X_{t})dW_{t}$ (5.4)
with $\mu(x)=((x_{1}-x_{1}^{3}-x_{2})/\alpha,\gamma x_{1}-x_{2}+\beta)^{\top}$ and $\sigma(x)=\sigma_{0}\mathtt{I}_{d_{x}}$. This is an elliptic diffusion with constant diffusion term. We fix the model parameters to $\alpha=0.1$, $\gamma=1$, $\beta=0.2$ and $\sigma_{0}=0.1$. We choose the starting values to be $X_{0}=(0.5,0.5)^{\top}$. We simulate the dynamics in the interval $t\in[0,10]$ using the Euler-Maruyama scheme with stepsize $\Xi=5\times 10^{-3}$ and observe the first component at regular times with spacing $10^{-2}$ and finally $\Delta=10^{-8}$ is the noise level. This scenario leads to 20 discretization points between two consecutive observations and 2000 in total.
We consider an auxiliary process with linearized drift $\tilde{\mu}(x)=\mu(x_{\text{ref}})+\frac{\partial}{\partial x}\mu(x_{\text{ref}})(x-x_{\text{ref}})$. We can calculate the reference point by evolving the true dynamics deterministically starting from $x$ using the stepsize $\Xi$ for 20 iterations until we reach the next observation time. The auxiliary process is an Ornstein-Uhlenbeck process with a transition density of $x^{\prime}$ given $x$ over time step $h$ as $\mathcal{N}_{2}(m,Q)$ where
$m=\exp(Jh)x+J^{-1}(\exp(Jh)-I)(\mu-Jx_{\text{ref}})$
$Q=\exp(Jh)\Psi\exp(J^{\top}h)-\Psi$
where $J=\frac{\partial}{\partial x}\mu(x_{\text{ref}})$, $\mu=\mu(x_{\text{ref}})$ and $\text{vec}(\Psi)=(I\otimes J+J\otimes I)^{-1}\text{vec}(\Sigma)$, $\otimes$ is a Kronecker product and $\exp(Jh)$ is a matrix exponentiation. Here, we assume that the jacobian $J$ is both diagonalizable and invertible.
We consider 4 PFs, the low noise and degenerate cases have been detailed. We also consider a ‘bootstrap’ and ‘optimal’ PF which are a little different from the conventional ones we have used so far. Both the bootstrap and optimal methods consider a smoother of the type (4.6) except under the natural parameterization. That is, after time discretization, the target smoother is up-to a constant at time $n$
$\prod_{k=1}^{n}\Big\{h_{k}(y_{k}|x_{k})R_{k-1,k}^{l}(C_{k-1,k}^{l}(x_{k-1},\mathbf{w}_{[k-1,k]}^{l},x_{k}))\mathbb{W}(d\mathbf{w}_{[k-1,k]}^{l})\lambda_{d_{x}}(x_{k})\Big\}.$
The bootstrap PF will use a proposal that is exactly the auxiliary process (on $x_{k}$ the Brownian motion is sampled from its law) whereas the optimal (which is not optimal in any sense) is a proposal proportional to the term $h_{k}(y_{k}|x_{k})$ multiplied by the density of the auxiliary process and the Brownian motion sampled from its law.
In Figure 8 we plot the ESS against time for a particular run. Here, we have set $N=10^{4}$. As in the previous example, the bootstrap PF collapses completely and the other filters perform quite well for this criterion. This shows that one can filter in low or degenerate noise and with a high amount of time discretization. This deduction is supported by the results in Figures 9 (marginal densities of the state) and 10 (tracking the true signal) which shows that one can reconstruct the signal process in this case. In Figure 11 we repeat the experiment for $\Delta\to 0$, where we see how the bootstrap PF collapses and the low-noise PF converges to the degenerate case.
Figure 8: ESS versus time. This is for the diffusion model.
Figure 9: Marginals of $X_{t,2}$ for $t\in\{0.1,3.4,6.7,10\}$. This is for the diffusion model.
Figure 10: Sample paths (top) and the mean estimates (bottom) versus time. This is for the diffusion model.
Figure 11: ESS versus $\log_{10}(1/\Delta)$. This is for the diffusion model.
5.3.1 State-dependent Diffusion Model
A more interesting example is when the diffusion term in FitzHugh-Nagumo model depends on the state. We now revisit the previous model with exactly the same setting, except with the diffusion coefficient
$\sigma(x)=\begin{bmatrix}\sigma_{1}\sqrt{x_{1}^{2}+1}&0\\ 0&\sigma_{2}\sqrt{x_{2}^{2}+1}\end{bmatrix}$ (5.5)
where we fix $\sigma_{1}=\sigma_{2}=0.1$. We remark that, contrary to the previous example (constant diffusion coefficient), we do not know of any potential alternative PF method. In the constant diffusion case, one could envisage using the Girsanov method w.r.t. a linear diffusion with the same constant drift, but that is not possible in this case. The auxiliary process is as before except with $\tilde{\sigma}=\sigma(x^{\prime})$, $x^{\prime}$ the end-point. To actually propose the new position of the particle instead of using the auxiliary process, we use an approximation of the auxiliary density by substituting $x^{\prime}$ with deterministic $x_{\text{ref}}$ and using $\sigma(x_{\text{ref}})$. This incurs an extra correction term in the importance weights as now one needs to divide by the associated density. The ‘bootstrap’ and ‘optimal’ methods use this new density in lieu of the auxiliary process density used in the constant diffusion case. We run each PF with $N=10^{4}$ particles.
We consider the ESS versus $\log_{10}(1/\Delta)$ in Figure 12. We observe similar results to the case of a constant diffusion coefficient. This shows that one can use the method for non-constant diffusions in low noise and high time discretization error.
Figure 12: ESS versus $\log_{10}(1/\Delta)$. This is for the state-dependent diffusion model.
Appendix A Proofs
A.1 Proof of Proposition 3.1
Proof.
As was recognized in [21] we have that for any fixed $\widetilde{z}\in\mathbb{R}^{d_{x}}$
$u_{n}^{\Delta}(\widetilde{z})\rightarrow\begin{bmatrix}x_{n}^{\star}\\ 0\end{bmatrix}+\begin{bmatrix}V_{n}z\\ \bar{z}\end{bmatrix}$. (A.1)
Therefore, by the dominated convergence theorem we have for any $\varphi\in\mathtt{B}_{b}(\mathbb{R}^{d_{x}-d_{y}})$
$\lim_{\Delta\downarrow 0}\int_{\mathbb{R}^{nd_{x}}}\varphi(z_{n})\prod_{k=1}^{n}p_{k}(u_{k}^{\Delta}(\widetilde{z}_{k},\epsilon))f_{k}(u_{k}^{\Delta}(\widetilde{z}_{k},x)|u_{k-1}^{\Delta}(\widetilde{z}_{k-1},x))\bigotimes_{i=1}^{n}\lambda_{d_{x}}(d\widetilde{z}_{i})=$
$\int_{\mathbb{R}^{nd_{x}}}\varphi(z_{n})\prod_{k=1}^{n}p_{k}(\overline{z}_{k})f_{k}(u_{k}^{\star}(z_{k})|u_{k-1}^{\star}(z_{k-1}))\bigotimes_{i=1}^{n}\lambda_{d_{x}}(d\widetilde{z}_{i})=$
$\int_{\mathbb{R}^{n(d_{x}-d_{y})}}\varphi(z_{n})\prod_{k=1}^{n}f_{k}(u_{k}^{\star}(z_{k})|u_{k-1}^{\star}(z_{k-1}))\bigotimes_{i=1}^{n}\lambda_{d_{x}}(dz_{i})$
and hence that
$\lim_{\Delta\downarrow 0}\pi_{n}^{\Delta}(\varphi)=\pi_{n}^{\star}(\varphi).$
Now, using (A.1) and assumption (A(A1)) we have for any fixed $(\widetilde{z}^{\prime},\widetilde{z})\in\mathbb{R}^{d_{x}}$
$\lim_{\Delta\downarrow 0}w_{n}^{\Delta}(\widetilde{z}^{\prime},\widetilde{z})=\frac{p_{n}(\overline{z})f_{n}(u_{n}^{\star}(z)|u_{n-1}^{\star}(z^{\prime}))}{p_{n}(\overline{z})q_{n}(z|z^{\prime})}=w_{n}^{\star}(z^{\prime},z).$ (A.2)
Now the joint density of the PF in the low noise case is
$\mathbb{P}^{\Delta}_{n}(\widetilde{z}_{1:n}^{1:N},\widetilde{a}_{1:n-1}^{1:N})=\left\{\prod_{i=1}^{N}\widetilde{q}_{1}^{\Delta}(\widetilde{z}_{1}^{i}|x_{0})\right\}\left\{\prod_{k=2}^{n}\prod_{i=1}^{N}\frac{w_{k-1}^{\Delta}(\widetilde{z}_{k-2}^{\widetilde{a}_{k-2}^{i}},\widetilde{z}_{k-1}^{\widetilde{a}_{k-1}^{i}})}{\sum_{s=1}^{N}w_{k-1}^{\Delta}(\widetilde{z}_{k-2}^{\widetilde{a}_{k-2}^{s}},\widetilde{z}_{k-1}^{s})}\widetilde{q}^{\Delta}(\widetilde{z}_{k}^{i}|\widetilde{z}_{k-1}^{\widetilde{a}_{k-1}^{i}})\right\}.$
Therefore, applying (A.2) and (A(A1)) for any fixed $\widetilde{z}_{1:n}^{1:N},\widetilde{a}_{1:n-1}^{1:N}$ we have that
$\lim_{\Delta\downarrow 0}\mathbb{P}^{\Delta}_{n}(\widetilde{z}_{1:n}^{1:N},\widetilde{a}_{1:n-1}^{1:N})=$
$\left\{\prod_{i=1}^{N}p_{1}(\overline{z}_{1}^{i})q_{1}^{\star}(z_{1}^{i}|x_{0})\right\}\left\{\prod_{k=2}^{n}\prod_{i=1}^{N}\frac{w_{k-1}^{\star}(z_{k-2}^{\widetilde{a}_{k-2}^{i}},z_{k-1}^{\widetilde{a}_{k-1}^{i}})}{\sum_{s=1}^{N}w_{k-1}^{\star}(z_{k-2}^{\widetilde{a}_{k-2}^{s}},z_{k-1}^{s})}p_{n}(\overline{z}_{k}^{i})q^{\star}(z_{k}^{i}|z_{k-1}^{\widetilde{a}_{k-1}^{i}})\right\}.$
So by the dominated convergence theorem we have that
$\lim_{\Delta\downarrow 0}\mathbb{E}^{\Delta}\left[\left|\pi_{n}^{\Delta,N}(\varphi)-\pi_{n}^{\Delta}(\varphi)\right|^{r}\right]=$
$\sum_{\widetilde{\mathbf{a}}_{1:n-1}^{1:N}\in\{1,\dots,N\}^{n-1}}\int_{\mathbb{R}^{Nnd_{x}}}\left|\frac{\sum_{i=1}^{N}\varphi(z_{n}^{i})w_{n}^{\star}(z_{n-1}^{\widetilde{a}_{n-1}^{i}},z_{n}^{i})}{\sum_{i=1}^{N}w_{n}^{\star}(z_{n-1}^{\widetilde{a}_{n-1}^{i}},z_{n}^{i})}-\pi_{n}^{\star}(\varphi)\right|^{r}\times$
$\left\{\prod_{i=1}^{N}p_{1}(\overline{z}_{1}^{i})q_{1}^{\star}(z_{1}^{i}|x_{0})\right\}\left\{\prod_{k=2}^{n}\prod_{i=1}^{N}\frac{w_{k-1}^{\star}(z_{k-2}^{\widetilde{a}_{k-2}^{i}},z_{k-1}^{\widetilde{a}_{k-1}^{i}})}{\sum_{s=1}^{N}w_{k-1}^{\star}(z_{k-2}^{\widetilde{a}_{k-2}^{s}},z_{k-1}^{s})}p_{n}(\overline{z}_{k}^{i})q^{\star}(z_{k}^{i}|z_{k-1}^{\widetilde{a}_{k-1}^{i}})\right\}\bigotimes_{k=1}^{n}\bigotimes_{i=1}^{N}\lambda_{d_{x}}(d\widetilde{z}^{i}_{k})=$
$\sum_{\widetilde{\mathbf{a}}_{1:n-1}^{1:N}\in\{1,\dots,N\}^{n-1}}\int_{\mathbb{R}^{Nn(d_{x}-d_{y})}}\left|\frac{\sum_{i=1}^{N}\varphi(z_{n}^{i})w_{n}^{\star}(z_{n-1}^{\widetilde{a}_{n-1}^{i}},z_{n}^{i})}{\sum_{i=1}^{N}w_{n}^{\star}(z_{n-1}^{\widetilde{a}_{n-1}^{i}},z_{n}^{i})}-\pi_{n}^{\star}(\varphi)\right|^{r}\times$
$\left\{\prod_{i=1}^{N}q_{1}^{\star}(z_{1}^{i}|x_{0})\right\}\left\{\prod_{k=2}^{n}\prod_{i=1}^{N}\frac{w_{k-1}^{\star}(z_{k-2}^{\widetilde{a}_{k-2}^{i}},z_{k-1}^{\widetilde{a}_{k-1}^{i}})}{\sum_{s=1}^{N}w_{k-1}^{\star}(z_{k-2}^{\widetilde{a}_{k-2}^{s}},z_{k-1}^{s})}q^{\star}(z_{k}^{i}|z_{k-1}^{\widetilde{a}_{k-1}^{i}})\right\}\bigotimes_{k=1}^{n}\bigotimes_{i=1}^{N}\lambda_{d_{x}}(dz^{i}_{k})$
where $\widetilde{\mathbf{a}}_{1:n-1}^{1:N}=(\widetilde{a}_{1}^{1},\dots,\widetilde{a}_{1}^{N},\dots,\widetilde{a}_{n-1}^{1},\dots,\widetilde{a}_{n-1}^{N})$. We recognize the last expression as the expectation w.r.t. the law of the degenerate noise PF; i.e. we have proved that
$\lim_{\Delta\downarrow 0}\mathbb{E}^{\Delta}\left[\left|\pi_{n}^{\Delta,N}(\varphi)-\pi_{n}^{\Delta}(\varphi)\right|^{r}\right]^{1/r}=\mathbb{E}^{\star}\left[\left|\pi_{n}^{\star,N}(\varphi)-\pi_{n}^{\star}(\varphi)\right|^{r}\right]^{1/r}.$
A.2 Proof of Proposition 3.2
Proof.
The results under the natural parameterization are simple consequences of the choice of the proposal. For the low-noise case using simple linear algebra one can verify that
$p_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},\epsilon))f_{n}(u_{n}^{\Delta}(\widetilde{z}_{n},x)|u_{n-1}^{\Delta}(\widetilde{z}_{n-1},x))\propto\exp\left\{-\frac{1}{2}z_{n}^{\top}\widetilde{\Omega}_{n}^{-1}z_{n}+\widetilde{\mu}_{n}^{\top}z_{n}\right\}.$
From there one has the optimal proposal and the normalizing constant, which is (up-to a constant) $w_{n}^{\Delta,\text{opt}}(\widetilde{z}_{n-1})$ can be obtained by completing the square and using Gaussian integrals. For the degenerate case we have similarly that
$f_{n}(u_{n}^{\star}(z_{n})|u_{n-1}^{\star}(z_{n-1}))\propto\exp\left\{-\frac{1}{2}z_{n}^{\top}(\Omega_{n}^{\star})^{-1}z_{n}+(\mu_{n}^{\star})^{\top}z_{n}\right\}$
and the proof is similar to the low-noise case. The final limit is a direct consequence of (A.1) and the dominated convergence theorem. ∎